If a car travels first 40km at speed of 20km/h and next 80km at a speed of 40km/h. what is the
average speed of car during journey?

Grade:9

1 Answers

Ruz
2 years ago
By definition, the average speed is given by-
 
\frac{total distance travelled}{total time taken}
 
We already know the total distance travelled, which is: 40km+ 80km= 120km
 
For calculating time, we use the equation s=v*t or t=s/v 
that is, time taken= distance/speed
 
For first part, time taken, t1= 40/20= 2 hours
For second part, time taken, t2= 80/40= 2 hours
 
So total time= 4 hours 
 
Putting this is our formula, avg speed= total distance/total time
We get, avg speed= 120/4= 30 km/h